2x+9times the sq.root of x=5
Answer by stanbon(75887)

2x+9sqrt(x) = 5
9sqrt(x) = 5-2x
Square both sides to get:
81x = 25-20x+4x^2
4x^2-101x+25=0
x = [101 +- sqrt(101^2-4*4*25)]/8
x = [101 +- sqrt(9801)]/8
x = [101 +- 99]/8
x = 1/4 or x = 25
===========
Cheers,
Stan H.